OpenStudy (anonymous):

solve x^2+10x=24 by completing the square.

OpenStudy (anonymous):

ok so first you would want to divide 10 by 2 , and you get 5. Now you want to square 5, and you get 25. so now you add 25 to the equation, and what you do to one side you do to the other x^2+10x+25=24+25 x^2+10x+25=49 but to solve, you equal it to zero. x^2+10x-24=0 now you would factor this. (x-2) (x+12) equal each of these to zero and solve x = 2 and x=-12

OpenStudy (anonymous):

Thank you idontsuck. and you don't suck you are really good in math! well at least better then me.

OpenStudy (anonymous):

aw thanks i really like you

OpenStudy (anonymous):

cool you are really nice.

OpenStudy (anonymous):

iDontSuck, do you midn looking at my multivariable calculus question?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

cool thats wonderful. i don't think of my self as a creep.